### v1

@@ Abstract
Point cloud place recognition (PCPR) plays a crucial role in photogrammetry and robotics applications such as autonomous driving, intelligent transportation, and augmented reality. In real-world large-scale deployments of a positioning system, PCPR models must continuously acquire, update, and accumulate knowledge to adapt to diverse and dynamic environments, i.e., the ability known as continual learning (CL). However, existing PCPR models often suffer from catastrophic forgetting, leading to significant performance degradation in previously learned scenes when adapting to new environments or sensor types. This results in poor model scalability, increased maintenance costs, and system deployment difficulties, undermining the practicality of PCPR. To address these issues, we propose LifelongPR, a novel continual learning framework for PCPR, which effectively extracts and fuses knowledge from sequential point cloud data. First, to alleviate the knowledge loss, we propose a replay sample selection method that dynamically allocates sample sizes according to each dataset’s information quantity and selects spatially diverse samples for maximal representativeness. Second, to handle domain shifts, we design a prompt learning-based CL framework with a lightweight prompt module and a two-stage training strategy, enabling domain-specific feature adaptation while minimizing forgetting. Comprehensive experiments on large-scale public and self-collected datasets are conducted to validate the effectiveness of the proposed method. Compared with the state-of-the-art (SOTA) method, our method achieves 6.50% improvement in $mIR@1$, 7.96% improvement in $mR@1$, and an 8.95% reduction in $F$. The code and pre-trained models are publicly available at https://github.com/zouxianghong/LifelongPR.
keywords:
[…]
{highlights}
Remarkable anti-forgetting capability for point cloud place recognition.
Reasonable replay sample selection based on datasets’ information quantity and samples’ spatial distribution.
Novel continual learning framework based on prompt learning.
1 Introduction
@@
Place recognition is a foundational task in photogrammetry and robotics, enabling autonomous systems to localize themselves within previously mapped environments by identifying revisited places (Li et al., 2021, 2025). It serves as a key module in a wide range of applications, including autonomous driving (Yurtsever et al., 2020; Li et al., 2023), intelligent transportation systems (Ounoughi and Yahia, 2023; Li et al., 2024), collaborative mapping (Tessier et al., 2010; Buehrer et al., 2018), and augmented reality (Carmigniani et al., 2011; Li et al., 2025). In recent years, point cloud place recognition (PCPR) has gained increasing attention due to its robustness against illumination changes, occlusions, and seasonal variations. With the advancement of deep learning techniques, PCPR has achieved remarkable progress in both recognition accuracy and robustness (Uy and Lee, 2018; Komorowski, 2021; Zou et al., 2023), making it a promising solution for long-term and all-weather localization in complex environments.
@@
Despite this progress, real-world large-scale deployments of PCPR introduce new challenges. PCPR models are often required to operate across diverse scenes (urban, suburban, or campus) and under different LiDAR sensor types as shown in Fig. 1. Point clouds collected under such heterogeneous conditions exhibit significant domain shifts, which cause severe performance degradation when the model is applied beyond its original training data (Jung et al., 2024). To ensure consistent performance, the PCPR model must be continually updated with new data collected over time and across domains (Lesort et al., 2020). However, such continual adaptation leads to catastrophic forgetting, where learning new environments overwrites knowledge of previously seen ones, undermining model reliability and safety. For instance, a model fine-tuned on a new city or sensor may lose up to 30% recall on earlier deployment zones (see column 1 in Fig. 4.2 and Fig. 4.2). A naive solution would be to retrain the model jointly with all data, which is computationally expensive and unscalable. Alternatively, maintaining separate models for each scenario drastically increases storage and system complexity (Aljundi et al., 2017; Serra et al., 2018). These limitations highlight the urgent need for effective knowledge fusion frameworks that can incrementally learn from new data while retaining prior knowledge, enabling scalable and robust PCPR in real-world, lifelong applications.
Fig. 1.: Catastrophic forgetting in the evolution process of localization models.
@@
Continual learning (CL) studies the problem of fusing knowledge from dynamic data distribution, aiming at mitigating catastrophic forgetting (Lesort et al., 2020; Zheng et al., 2025). In recent years, a few studies have introduced the theory of CL to mitigate catastrophic forgetting in PCPR (Knights et al., 2022; Cui and Chen, 2023). These methods primarily rely on two strategies: sample replay and knowledge distillation. Replay-based methods store a limited number of past samples and interleave them with new data during training, while distillation-based approaches transfer knowledge from previous models by minimizing output divergence(Rebuffi et al., 2017; Li and Hoiem, 2017). While these approaches offer partial mitigation, they suffer from two fundamental limitations that significantly constrain their effectiveness in real-world deployments. First, existing methods typically adopt random sampling for replay memory, ignoring the information content and spatial distribution of training data. This leads to unrepresentative replay sets that fail to preserve essential scene knowledge, especially in large and diverse environments. Second, when domain gaps between training stages are substantial (e.g., due to changes in environment or LiDAR sensor), a small replay buffer becomes insufficient to retain old knowledge, while the effectiveness of knowledge distillation also declines due to increased feature drift.
@@
The remainder of this article is organized as follows. Section 2 reviews related works. Section 3 details the proposed method. Section 4 presents experimental data, and quantitative evaluation. Section 5 presents the ablation studies and discusses the deficiencies and future work. Finally, Section 6 concludes the paper.
2 Related Work
2.1 Learning-based point cloud place recognition
@@
Projection-based methods extract features from 2D images (e.g., range images and bird’s eye view (BEV) images) generated from raw point clouds. OREOS (Schaupp et al., 2019) feeds range images generated from point clouds into a CNN network and extracts features for characterizing position and orientation respectively, enabling simultaneous place recognition and yaw estimation. Kim et al. (2019) generates BEV images from point clouds based on Scan Context (Kim and Kim, 2018) and regards the PCPR as a classification task. OverlapNet (Chen et al., 2021) feeds range images to a Siamese neural network and directly regresses the overlap ratio and yaw angle of two point clouds. While computationally efficient, these methods incur significant information loss in the process of converting point clouds into 2D images.
Voxel-based methods extract features from voxelized point clouds. Minkloc3D (Komorowski, 2021) proposes the first voxel-based PCPR method, extracting local and contextual features via feature pyramid and sparse convolution, while aggregating global features with GeM. Komorowski (2022) enhances it with the ECA attention (Wang et al., 2020) module and average precision loss. SVT-Net (Fan et al., 2022) combines the sparse convolution and transformer to extract global features, making the model lightweight and efficient. Such methods have simple network structures and are more noise resistant, while voxelization suffers from information loss.
Point-based methods extract features directly from raw point clouds. Uy and Lee (2018) propose the first learning-based PCPR method built upon PointNet (Qi et al., 2017) and NetVLAD (Arandjelovic et al., 2016). Hui et al. (2021) utilizes the EdgeConv and UNet-like structure to enhance the feature representation ability of the network. Zhou et al. (2021) feeds probabilistic representation converted from point clouds into the network, increasing the noise immunity. PatchAugNet (Zou et al., 2023) enhances the adaptability of the network to domain gaps between heterogeneous point clouds with a patch feature augmentation module and an adaptive pyramid feature aggregation module. LAWS (Xie et al., 2024) treats place recognition as a classification task and eliminates classification ambiguity through two spatial partitioning patterns. Such methods retain the information of original point clouds, but they have poor generalization ability to unseen scenes.
@@ 2.2 Continual learning
CL, also termed incremental learning or lifelong learning, learns from dynamic data distributions, and its key problem is the trade-off between learning plasticity and memory stability (Wang et al., 2024). Existing CL methods primarily fall into three categories: replay-based, regularization-based, and parameter isolation-based (De Lange et al., 2021).
@@
Replay-based methods first keep original samples from previous tasks or generate pseudo samples from a generative model and then learn jointly with current task data. The most representative method is iCaRL (Rebuffi et al., 2017). Aiming at class incremental learning tasks, it selects replay samples according to the distance between replay samples and class centers and applies knowledge distillation to mitigate catastrophic forgetting. BIC (Wu et al., 2019) utilizes class-balanced validation sets to train deviation parameters for network output adjustment, addressing the problem of imbalanced new and replay sample sizes in class incremental learning tasks. DGR (Shin et al., 2017), the first CL method utilizing the generative model, alternately trains the task network and generative network. MeRGAN (Wu et al., 2018) further utilizes replay alignment to enhance the consistency between old and new generative models, playing a role similar to regularization. Such methods usually mitigate catastrophic forgetting well, and its key point lies in replay sample selection and utilization (MacQueen, 1967; Killamsetty et al., 2021; Kothawade et al., 2022).
@@
Regularization-based methods balance current and previous tasks through explicit regularization items, falling into two categories: weight regularization-based and function regularization-based. Weight regularization-based methods selectively regularize the variation of network parameters by evaluating the importance of network parameters to previous tasks (Kirkpatrick et al., 2017). EWC (Kirkpatrick et al., 2017) is the most representative weight regularization-based method. Function regularization-based methods utilize the previous models to guide the training of the current model through knowledge distillation. LwF (Li and Hoiem, 2017) first introduces knowledge distillation to CL and regards the output of previous models as soft labels for training the current model. Such methods are simple and effective but sensitive to domain gaps between tasks.
@@
Parameter isolation-based methods allocate dedicated parameters for specific tasks and can guarantee maximal stability by fixing the parameter subsets of previous tasks. Methods like HAT (Serra et al., 2018) and PackNet (Mallya and Lazebnik, 2018) retain multiple task-specific models with identical network structures. Other methods like Expert Gate (Aljundi et al., 2017) and PathNet (Fernando et al., 2017) divide the network into task-sharing and task-specific parts and limit the interference between tasks to the task-specific network structure. Such methods avoid or mitigate catastrophic forgetting via task-specific networks, while they must maintain multiple models.
@@
In recent years, the above CL methods have been introduced to PCPR for catastrophic forgetting mitigation. InCloud (Knights et al., 2022), the first CL method for PCPR, utilizes sample replay and knowledge distillation to mitigate catastrophic forgetting. CCL (Cui and Chen, 2023) combines contrastive learning and CL to improve the network’s generalizability and mitigate catastrophic forgetting. Both methods’ random replay sampling fails to adequately represent historical training data. BioSLAM (Yin et al., 2023) achieves CL for cross-modal place recognition via GAN-based sample generation and dual-memory replay selection, but incurs increased training complexity and reduced efficiency.
2.3 Prompt learning for continual learning
@@
Prompt learning is an emerging technique, designing task-specific prompts to conditionally adapt pre-trained models for downstream tasks (Liu et al., 2023). Unlike Adapter (Houlsby et al., 2019) and LoRA (Hu et al., 2021), it avoids architectural modifications with minimal parameter additions. Some methods strike a balance between pre-trained knowledge and downstream tasks with lightweight prompts, serving as external memory for CL (Zhou et al., 2024). L2P (Wang et al., 2022a) is the first method to introduce prompt learning to CL. It proposes a learnable prompt pool to instruct the pre-trained model to perform downstream tasks conditionally. DualPrompt (Wang et al., 2022b) proposes dual prompts for CL. G-prompt and E-prompt learn task-invariant and task-specific instructions respectively. S-Prompt (Wang et al., 2022c) builds task centers by K-means clustering and utilizes the KNN search to find the most similar prompt. DAP (Jung et al., 2023) utilizes a lightweight MLP network to construct instance-level prompts to resolve domain scalability issues. Existing methods are predominantly designed for transformer-based networks, and how to use prompts for non-transformer networks still need further research.
3 Methodology
@@
Fig. 2.: Overview of the proposed method.
3.1 Definition of replay-based CL for PCPR
As shown in Fig. 3.1, replay-based CL for PCPR is a domain-incremental learning task, which uses $T$ discontinuous training sets $D_{1:T}=\{D_{t}\}^{T}_{t=1}$ to update the PCPR model $\mathrm{F}(\theta)$  [^1] The following section omits the symbol $\theta$. incrementally, aiming to achieve better place recognition performance in their respective scenes. Each training set $D_{t}$ is collected from different scenes or using different LiDAR sensors, leading to domain gaps among various training sets. When training on $D_{t}$, the complete historical training sets $D_{1:t-1}$ cannot be fully accessed; instead, only a subset $M_{1:t-1}\subset D_{1:t-1}$ can be retrieved for replay. Given constraints on training efficiency and limited storage resources, the total number of replay samples $k_{total}$ is typically fixed. After updating the place recognition model with the current training set and the replay samples $\hat{D}_{t}=D_{t}\cup M_{1:t-1}$, the model’s performance still tends to decline in the scenes corresponding to the historical training sets $D_{1:t-1}$, i.e. catastrophic forgetting. This is the core challenge in CL for PCPR.
Fig. 3.: Replay-based continual learning for point cloud place recognition.
3.2 Replay sample selection considering information quantity and spatial distribution
@@
3.2.1 Sample size allocation based on information quantity
Existing CL methods for PCPR almost uniformly select the same number of samples from all training sets for replay, overlooking variations in information quantity across training sets. The information quantity of a training set is related to factors such as the size and diversity of samples, and it offers clues as to how many samples to select from each training set. To this end, we introduce the rank of the kernel matrix (Fine and Scheinberg, 2001) to quantitatively assess the information quantity of training sets and, dynamically allocating the number of replay samples to select from each training set.

### v2

@@ Abstract
Point cloud place recognition (PCPR) determines the geo-location within a prebuilt map and plays a crucial role in geoscience and robotics applications such as autonomous driving, intelligent transportation, and augmented reality. In real-world large-scale deployments of a geographic positioning system, PCPR models must continuously acquire, update, and accumulate knowledge to adapt to diverse and dynamic environments, i.e., the ability known as continual learning (CL). However, existing PCPR models often suffer from catastrophic forgetting, leading to significant performance degradation in previously learned scenes when adapting to new environments or sensor types. This results in poor model scalability, increased maintenance costs, and system deployment difficulties, undermining the practicality of PCPR. To address these issues, we propose LifelongPR, a novel continual learning framework for PCPR, which effectively extracts and fuses knowledge from sequential point cloud data. First, to alleviate the knowledge loss, we propose a replay sample selection method that dynamically allocates sample sizes according to each dataset’s information quantity and selects spatially diverse samples for maximal representativeness. Second, to handle domain shifts, we design a prompt learning-based CL framework with a lightweight prompt module and a two-stage training strategy, enabling domain-specific feature adaptation while minimizing forgetting. Comprehensive experiments on large-scale public and self-collected datasets are conducted to validate the effectiveness of the proposed method. Compared with the state-of-the-art (SOTA) method, our method achieves 6.50% improvement in $mIR@1$, 7.96% improvement in $mR@1$, and an 8.95% reduction in $F$. The code and pre-trained models are publicly available at https://zouxianghong.github.io/LifelongPR.
Index Terms:
Point Cloud Place Recognition; Continual Learning; Catastrophic Forgetting; Replay Sample Selection; Prompt Learning
I Introduction
@@
Place recognition is a foundational task in geoscience and robotics, enabling autonomous systems to determining their geo-locations within previously mapped environments by identifying revisited places [1, 2]. It serves as a key module in a wide range of applications, including autonomous driving [3, 4], intelligent transportation systems [5, 6], collaborative mapping [7, 8], and augmented reality [9, 10]. In recent years, point cloud place recognition (PCPR) has gained increasing attention due to its robustness against illumination changes, occlusions, and seasonal variations. With the advancement of deep learning techniques, PCPR has achieved remarkable progress in both recognition accuracy and robustness [11, 12, 13], making it a promising solution for long-term and all-weather localization in complex environments.
@@
Despite this progress, real-world large-scale deployments of PCPR introduce new challenges. PCPR models are often required to operate across diverse scenes (urban, suburban, or campus) and under different LiDAR sensor types as shown in Fig. 1. Point clouds collected under such heterogeneous conditions exhibit significant domain shifts, which cause severe performance degradation when the model is applied beyond its original training data [14]. To ensure consistent performance, the PCPR model must be continually updated with new data collected over time and across domains [15]. However, such continual adaptation leads to catastrophic forgetting, where learning new environments overwrites knowledge of previously seen ones, undermining model reliability and safety. For instance, a model fine-tuned on a new city or sensor may lose up to 30% recall on earlier deployment zones (see column 1 in Fig. 6 and Fig. 7). A naive solution would be to retrain the model jointly with all data, which is computationally expensive and unscalable. Alternatively, maintaining separate models for each scenario drastically increases storage and system complexity [16, 17]. These limitations highlight the urgent need for effective knowledge fusion frameworks that can incrementally learn from new data while retaining prior knowledge, enabling scalable and robust PCPR in real-world, lifelong applications.
Figure 1: Lifelong point cloud place recognition encounters challenges posed by diverse localization scenes and LiDAR sensors. Above information mainly comes from the experimental datasets in Section IV.
@@
Continual learning (CL) studies the problem of fusing knowledge from dynamic data distribution, aiming at mitigating catastrophic forgetting [15, 18]. In recent years, a few studies have introduced the theory of CL to mitigate catastrophic forgetting in PCPR [19, 20]. These methods primarily rely on two strategies: sample replay and knowledge distillation. Replay-based methods store a limited number of past samples and interleave them with new data during training, while distillation-based approaches transfer knowledge from previous models by minimizing output divergence[21, 22]. While these approaches offer partial mitigation, they suffer from two fundamental limitations that significantly constrain their effectiveness in real-world deployments. First, existing methods typically adopt random sampling for replay memory, ignoring the information content and spatial distribution of training data. This leads to unrepresentative replay sets that fail to preserve essential scene knowledge, especially in large and diverse environments. Second, when domain gaps between training stages are substantial (e.g., due to changes in environment or LiDAR sensor), a small replay buffer becomes insufficient to retain old knowledge, while the effectiveness of knowledge distillation also declines due to increased feature drift.
@@
The remainder of this article is organized as follows. Section II reviews related works. Section III details the proposed method. Section IV presents experimental data, and quantitative evaluation. Section V presents the ablation studies and discusses the deficiencies and future work. Finally, Section VI concludes the paper.
II Related Work
II-A Learning-based point cloud place recognition
@@
Projection-based methods extract features from 2D images (e.g., range images and bird’s eye view (BEV) images) generated from raw point clouds. OREOS [23] feeds range images generated from point clouds into a CNN network and extracts features for characterizing position and orientation respectively, enabling simultaneous place recognition and yaw estimation. Kim et al. [24] generates BEV images from point clouds based on Scan Context [25] and regards the PCPR as a classification task. OverlapNet [26] feeds range images to a Siamese neural network and directly regresses the overlap ratio and yaw angle of two point clouds. While computationally efficient, these methods incur significant information loss in the process of converting point clouds into 2D images.
Voxel-based methods extract features from voxelized point clouds. Minkloc3D [12] proposes the first voxel-based PCPR method, extracting local and contextual features via feature pyramid and sparse convolution, while aggregating global features with GeM. Komorowski [27] enhances it with the ECA attention [28] module and average precision loss. SVT-Net [29] combines the sparse convolution and transformer to extract global features, making the model lightweight and efficient. Such methods have simple network structures and are more noise resistant, while voxelization suffers from information loss.
Point-based methods extract features directly from raw point clouds. Uy and Lee [11] propose the first learning-based PCPR method built upon PointNet [30] and NetVLAD [31]. Hui et al. [32] utilizes the EdgeConv and UNet-like structure to enhance the feature representation ability of the network. Zhou et al. [33] feeds probabilistic representation converted from point clouds into the network, increasing the noise immunity. PatchAugNet [13] enhances the adaptability of the network to domain gaps between heterogeneous point clouds with a patch feature augmentation module and an adaptive pyramid feature aggregation module. LAWS [34] treats place recognition as a classification task and eliminates classification ambiguity through two spatial partitioning patterns. Such methods retain the information of original point clouds, but they have poor generalization ability to unseen scenes.
@@ II-B Continual learning
CL, also termed incremental learning or lifelong learning, learns from dynamic data distributions, and its key problem is the trade-off between learning plasticity and memory stability [35]. Existing CL methods primarily fall into three categories: replay-based, regularization-based, and parameter isolation-based [36].
@@
Replay-based methods first keep original samples from previous tasks or generate pseudo samples from a generative model and then learn jointly with current task data. The most representative method is iCaRL [21]. Aiming at class incremental learning tasks, it selects replay samples according to the distance between replay samples and class centers and applies knowledge distillation to mitigate catastrophic forgetting. BIC [37] utilizes class-balanced validation sets to train deviation parameters for network output adjustment, addressing the problem of imbalanced new and replay sample sizes in class incremental learning tasks. DGR [38], the first CL method utilizing the generative model, alternately trains the task network and generative network. MeRGAN [39] further utilizes replay alignment to enhance the consistency between old and new generative models, playing a role similar to regularization. Such methods usually mitigate catastrophic forgetting well, and its key point lies in replay sample selection and utilization [40, 41, 42].
@@
Regularization-based methods balance current and previous tasks through explicit regularization items, falling into two categories: weight regularization-based and function regularization-based. Weight regularization-based methods selectively regularize the variation of network parameters by evaluating the importance of network parameters to previous tasks [43]. EWC [43] is the most representative weight regularization-based method. Function regularization-based methods utilize the previous models to guide the training of the current model through knowledge distillation. LwF [22] first introduces knowledge distillation to CL and regards the output of previous models as soft labels for training the current model. Such methods are simple and effective but sensitive to domain gaps between tasks.
@@
Parameter isolation-based methods allocate dedicated parameters for specific tasks and can guarantee maximal stability by fixing the parameter subsets of previous tasks. Methods like HAT [17] and PackNet [44] retain multiple task-specific models with identical network structures. Other methods like Expert Gate [16] and PathNet [45] divide the network into task-sharing and task-specific parts and limit the interference between tasks to the task-specific network structure. Such methods avoid or mitigate catastrophic forgetting via task-specific networks, while they must maintain multiple models.
@@
In recent years, the above CL methods have been introduced to PCPR for catastrophic forgetting mitigation. InCloud [19], the first CL method for PCPR, utilizes sample replay and knowledge distillation to mitigate catastrophic forgetting. CCL [20] combines contrastive learning and CL to improve the network’s generalizability and mitigate catastrophic forgetting. Both methods’ random replay sampling fails to adequately represent historical training data. BioSLAM [46] achieves CL for cross-modal place recognition via GAN-based sample generation and dual-memory replay selection, but incurs increased training complexity and reduced efficiency.
II-C Prompt learning for continual learning
@@
Prompt learning is an emerging technique, designing task-specific prompts to conditionally adapt pre-trained models for downstream tasks [47]. Unlike Adapter [48] and LoRA [49], it avoids architectural modifications with minimal parameter additions. Some methods strike a balance between pre-trained knowledge and downstream tasks with lightweight prompts, serving as external memory for CL [50]. L2P [51] is the first method to introduce prompt learning to CL. It proposes a learnable prompt pool to instruct the pre-trained model to perform downstream tasks conditionally. DualPrompt [52] proposes dual prompts for CL. G-prompt and E-prompt learn task-invariant and task-specific instructions respectively. S-Prompt [53] builds task centers by K-means clustering and utilizes the KNN search to find the most similar prompt. DAP [54] utilizes a lightweight MLP network to construct instance-level prompts to resolve domain scalability issues. Existing methods are predominantly designed for transformer-based networks, and how to use prompts for non-transformer networks still need further research.
III Methodology
@@
Figure 2: Overview of the proposed method. (a) overall workflow, (b) replay sample selection, (c) prompt module.
III-A Definition of replay-based CL for PCPR
As shown in Fig. 3, replay-based CL for PCPR is a domain-incremental learning task, which uses $T$ discontinuous training sets $D_{1:T}=\{D_{t}\}^{T}_{t=1}$ to update the PCPR model $\mathrm{F}(\theta)$  [^1] The following section omits the symbol $\theta$. incrementally, aiming to achieve better place recognition performance in their respective scenes. Each training set $D_{t}$ is collected from different scenes or using different LiDAR sensors, leading to domain gaps among various training sets. When training on $D_{t}$, the complete historical training sets $D_{1:t-1}$ cannot be fully accessed; instead, only a subset $M_{1:t-1}\subset D_{1:t-1}$ can be retrieved for replay. Given constraints on training efficiency and limited storage resources, the total number of replay samples $k_{total}$ is typically fixed. After updating the place recognition model with the current training set and the replay samples $\hat{D}_{t}=D_{t}\cup M_{1:t-1}$, the model’s performance still tends to decline in the scenes corresponding to the historical training sets $D_{1:t-1}$, i.e. catastrophic forgetting. This is the core challenge in CL for PCPR.
Figure 3: Replay-based continual learning for point cloud place recognition.
III-B Replay sample selection considering information quantity and spatial distribution
@@
III-B1 Sample size allocation based on information quantity
Existing CL methods for PCPR almost uniformly select the same number of samples from all training sets for replay, overlooking variations in information quantity across training sets. To this end, we quantitatively assess the information quantity of training sets and, dynamically allocating the number of replay samples to select from each training set.
Remark 1.
The information quantity of a training set is related to factors such as the size and diversity of samples, providing guidance on how many samples to select from each training set. The rank of the kernel matrix [55] represents the maximum number of linearly independent samples in the training set and thus reflects the magnitude of its information content, enabling quantitative assessment of the training set’s information quantity.